OpenStudy (automaticloveletter):

Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations? x + 1 = 4x + 5 x = 4x + 5 x + 1 = 4x x + 5 = 4x + 1

Mehek (mehek14):

which one do you think? you substitute the the first equation into the second equation

Mehek (mehek14):

in the first equation, y = x + 1 substitute that value for y in the second equation tell me what you get @automaticloveletter

OpenStudy (automaticloveletter):

5 = y-4x? (I'm using a calculator) :/

Mehek (mehek14):

no need for a calculator

Mehek (mehek14):

look the value of y in the first equation is x + 1 correct?

OpenStudy (automaticloveletter):

yes

Mehek (mehek14):

so when you substitute it into the second equation, you put x + 1 instead of y

OpenStudy (automaticloveletter):

wait so it's just 4*1+5?

Mehek (mehek14):

no you're not getting it.... y = x + 1 in the first equation second equation y = 4x + 5 put x + 1 instead of y in the second equation you get: x + 1 = 4x + 5

OpenStudy (automaticloveletter):

ohhh

Mehek (mehek14):

you just substitute the value in the first equation into the second they're not asking you to solve it

OpenStudy (automaticloveletter):

I know that, but I'd also want to know how to solve it for other questions..

Mehek (mehek14):

How you would solve for it: \(x + 1 = 4x + 5\\1 = 3x + 5\\-4=3x\\ -\dfrac{4}{3}=x\)

Mehek (mehek14):

@automaticloveletter